OpenStudy (anonymous):

PLEASE HELP WITH #3 !!!

OpenStudy (anonymous):

look at the numbers on the side, #3 please?!

OpenStudy (anonymous):

well you would use the equation 7.28x+8.40y=7.95

OpenStudy (anonymous):

isnt the equatoin 7.28x+8.40y=6(7.95) ? @sodukogirl13

OpenStudy (anonymous):

yes does the equation not help you out at all

OpenStudy (anonymous):

first figure out either they y or the x

OpenStudy (anonymous):

I meant to say first figure out either the y or the x

OpenStudy (anonymous):

to help I will figure out the y for you 7.28x+8.40y=6(7.95) 7.28x+8.40y=47.7 -7.28x -7.28x 8.40y=47.7-7.28x divide both sides by 8.40 y=5.68-0.87x now just plug in the left side for y in the original equation

OpenStudy (anonymous):

does that help you at all